under what conditions it has been following matrix equality?
( x y) * (1 1) = (1 1) * (x y)
( z w) (0 1) .... (0 1) (z w)
Answer: provided that z = 0 and x = w
How? Thank you

Answer by stanbon(75887)

under what conditions it has been following matrix equality?
( x y) * (1 1) = (1 1) * (x y)
( z w) * (0 1) = (0 1) * (z w)
Answer: provided that z = 0 and x = w
How?
---
Equations:
x + 0*y = x + z
x + y = y + w
z + 0*w = z + w
z + w = 0y + w
--------------------
From the 2nd row: x = w
From the 4th row: z = 0
===============================
Cheers,
Stan H.